OpenStudy (anonymous):

which equation represents a line that is parallelto the line represented by the equation y=3/5x+5/3? _ y=5/3x+5/3 _ y=3/5x+7 _ y=5/3x+3/5 _y=-3/5x+1/7

OpenStudy (anonymous):

-y=-3/5x+1/7 because in order for a line to be parallel they must have the same slope. In this case each slope will be m=3/5. Simplifying this equation would be dividing each side by -1 which would get you y=3/5x+1/7. Therefore they have the same slope.